OpenStudy (anonymous):

Find the slope of the line through the points (-1, -3) and (-1, 7)

OpenStudy (saifoo.khan):

7+3 ----- -1 +1

OpenStudy (anonymous):

The slope is undefined

OpenStudy (saifoo.khan):

zero. + undefined.

OpenStudy (anonymous):

10/0 :}

OpenStudy (anonymous):

\[m = (y2 - y1) / (x2 - x1)\] m = ( 7 -(-3)) / (-1-(-1)) = 10/ 0 = undefinited